OpenStudy (anonymous):

I know that this equation should equal either -2.162 or 4.162, but whenever I try to solve it, I always come up with the exact opposite: 2.162 or -4.162. Can someone help walk me through this? The equation is: 0=-x^2+2x+9

OpenStudy (anonymous):

what is a, b, c on the expression??

OpenStudy (anonymous):

a=-1 b=2 c=9 @OOOPS

OpenStudy (anonymous):

Right now use the quadratic formula \(\large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) Now plug them in:\(\large x=\frac{-2\pm\sqrt{(2)^2-4(-1)(9)}}{2(-1)}\) Did you get this far?

OpenStudy (aum):

\( x = \Large \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a } = \frac{ -2 \pm \sqrt{2 ^{2} - 4(-1)(9)} }{ 2(-1) = } = \\ \Large \frac{ -2 \pm 2\sqrt{10)} }{ -2 } = 1 \mp \sqrt{10} = ? \)

OpenStudy (anonymous):

I got it! Thank you! @venny and @aum